OpenStudy (anonymous):

Write the given expression as a single logarithm: 4log(5x) - 5[log(x) - log(y-5)]

OpenStudy (anonymous):

Can you write this one in single logarithm? log(x) - log(y-5)

OpenStudy (anonymous):

I don't know how to

OpenStudy (anonymous):

log a - logb = log (a/b) Can you do it now?

OpenStudy (anonymous):

So would it be log(x/y-5)?

OpenStudy (anonymous):

Yep, to be exact: log ( x / ( y -5) )

OpenStudy (anonymous):

so 5[log(x) - log(y-5)] = log ( x / ( y -5) )^5

OpenStudy (anonymous):

Now, do you know what to do with 4log(5x) ?

OpenStudy (anonymous):

No I was trying to figure it out but cant

OpenStudy (anonymous):

Did you see what I've done with number 5, then try to do similarly with number 4 : 4log(5x)

jhonyy9 (jhonyy9):

4log(5x) -5(log(x) - log(y-5)) = log(5x)^4 -5(log(x/(y-5))) = = log(5x)^4 - log(x/(y-5))^5 =log((625x^4)/((x/(y-5))^5) = = log(625x^4)/((x^5)/(y-5)^5) = log(((625x^4)(y-5)^5)/(x^5)) = = log((625(y-5)^5)/x)